) Matthew is an addicted coffee drinker and proud patron of Starbucks, so he keeps an eye out on the prices of coffee. He finds out that Starbucks increases its price of a grande frappuccino from $3.50 to $4.00, so he expects many patrons to consume less grande lattes from 2 to 1 per week, and to find an alternative. He lives by a Coffee Bean and Tea Leaf café and his opportunity for a grande frappucino is a medium ice blended. He plans to consume more medium ice blended drinks increasing his quantity demanded from 1 to 5 per week. What is the cross price elasticity of demand for a medium ice blended drink with respect to a grande frappuccino’s price?
1
Expert's answer
2020-05-13T11:09:21-0400
"p_1=3.5, p_2=4"

"Q_1=1, Q_2=5"


"E=\frac {Q_1-Q_2}{Q_1}\times \frac{p_1}{p_1-p_2}"


"E=\frac {1-5}{1}\times \frac{3.5}{3.5-4}=28"